### 26 Jan 2024 v1

# Attitudes Towards and Knowledge of Non-Consensual Synthetic Intimate Imagery in 10 Countries

Abstract.
Deepfake technology tools have become ubiquitous, ”democratizing” the ability to manipulate images and videos. One popular use of such technology is the creation of sexually explicit content, which can then be posted and shared widely on the internet. This article examines attitudes and behaviors related to non-consensual synthetic intimate imagery (NSII) across over 16,000 respondents in 10 countries. Despite nascent societal awareness of NSII, NSII behaviors were considered harmful. In regards to prevalence, 2.2% of all respondents indicated personal victimization, and 1.8% all of respondents indicated perpetration behaviors. Respondents from countries with relevant legislation also reported perpetration and victimization experiences, suggesting legislative action alone is not a sufficient solution to deter perpetration. Technical considerations to reduce harms may include suggestions for how individuals can better monitor their presence online, as well as enforced platform policies which ban, or allow for removal of, NSII content.
deepfake, pornography, image-based abuse, non-consensual explicit imagery, involuntary synthetic pornographic imagery
copyright: none
1. Introduction
Digitally altered content is a long-standing phenomenon, traditionally created through the use of photo-editing software to stitch together segments, add filters, or change the speed of videos. These older forms have now been dubbed ”cheapfakes” in contrast to ”deepfakes,” a portmanteau of ”deep learning” and ”fake” that describes AI-generated highly realistic content (Shahid et al., 2022; Westerlund, 2019). While deepfakes can be used in beneficial ways for accessibility and creativity (Farid, 2022; Cover, 2022), deepfakes can also be weaponized and used for malicious purposes, including financial fraud, disinformation dissemination, and sexual extortion (”sextortion”) (Farid, 2022; Aliman et al., 2021). Deepfake videos can be categorized into a variety of different types, including audio deepfakes, lip-syncing, puppet-master, face synthesis and attribute manipulation, and face-swapping (Wodajo and Atnafu, 2021). The focus of this paper is on face-swapping in the creation of deepfake pornography. This involves the face of the source person being replaced with the target person, so that the target person appears to engage in scenarios in which they never appeared (Masood et al., 2023). The terminology and language used has evolved in the image-based sexual abuse space more broadly (McGlynn et al., 2017) and accordingly, it is important to emphasize that deepfake pornography is, by its nature, involuntary and non-consensual. Therefore, throughout the paper, we use the term ”non-consensual synthetic intimate imagery” (NSII). One exception is when describing the language we used in the survey (where we used the more recognized ”deepfake pornography” term).
The consumer creation of deepfakes started in late 2017 on Reddit, after a user named ”deepfakes” posted NSII depicting the faces of female celebrities superimposed onto pornographic videos on his page (Kugler and Pace, 2021; Westerlund, 2019). Continued consumer interest in NSII is reflected in the proliferation of dedicated NSII sites, often depicting celebrity targets. Abuse potential has increased in recent years as the technology has advanced in sophistication and availability (Cao et al., 2023; Masood et al., 2023; Gieseke, 2020; Wang, 2022). NSII can be considered a form of image-based sexual abuse because intimate images are created and/or shared without the consent of the person or persons depicted in the images. The harms of image-based sexual abuse have been well-documented, including negative impacts on victim-survivors’ mental health, career prospects, and willingness to engage with others both online and offline (Henry et al., 2020; Citron and Franks, 2014). The proliferation of NSII technologies means that anyone can now become a victim of image-based sexual abuse, and research suggests that women continue to bear the brunt of this online abuse (Kugler and Pace, 2021; Dunn, 2021). Investigating the prevalence of the misuse of deepfake technologies, specifically for NSII, can help Human Computer Interaction (HCI) research communities to better understand how best to mitigate the gendered societal harms associated with this phenomenon. Potential avenues for reduced harm may include comprehensive and considered legislation, NSII policies around the creation and/or distribution of content (e.g., Reddit’s continued de-platforming of subreddits dedicated to deepfake pornography or Google’s policy for removal of NSII), technical tools to help victim-survivors discover and automate takedown requests, and user experience treatments on deepfake creation tools designed to deter users from creating and distributing NSII.
To begin mapping the prevalence of and general sentiment around NSII, we addressed the following research questions:
  • RQ1: What is the general public’s awareness of, and attitudes towards, NSII?
  • RQ2: What is the prevalence of NSII behaviors (e.g., creating, viewing, and or/sharing)?
  • RQ3: How does gender influence both RQ1 and RQ2?
In this study, we surveyed over 16,000 respondents in 10 different countries. Countries were selected based on a number of considerations, including representation of the diverse approach to legislation on the issue of image-based sexual abuse, replication/expansion of previous findings (e.g., in Australia and the United States) around the prevalence of image-based sexual abuse, and geographic diversity. Questions were designed to assess attitudes towards, and experiences of, image-based sexual abuse, with specific questions relating to NSII. This study was primarily quantitative, although optional open-ended questions allowed respondents to add additional context and thoughts. This research contributes to the existing literature by identifying trends in geographical awareness and behaviors; building on the existing image-based sexual abuse literature in terms of understanding who is most at-risk to this phenomenon; and probing the efficacy of current or proposed remedies.
2. Related Work
In this section, we provide a brief summary of the literature surrounding public perceptions of deepfakes more broadly and in relation to associated harms, deepfake detection, and legislative redress. The structure of this background review loosely follows that of our research questions; however, as prevalence data is a current gap in the literature, in our review below we also focus on whether existing or proposed laws sufficiently address the harms associated with this phenomenon.
2.1. Attitudes Towards Deepfakes
The implications of NSII both resemble, and are distinct to, traditional forms of image-based sexual abuse. In introducing the ”pervert’s dilemma,” Ohman (Öhman, 2020) notes that the mere creation of deepfake pornography in isolation is somewhat comparable to private sexual fantasies, which are not considered morally objectionable. However, when situated appropriately against the backdrop of gender inequality and gender-based violence, the moral unacceptability of NSII becomes clear. Moreover, deepfake technology provides a unique avenue for harassment and harm as it significantly reduces the barriers to the creation of non-consensual content (Akerley, 2020). As Harris (Harris, 2021) notes, while some have argued that concerns about the epistemic effects of deepfakes are overblown (e.g., because content is unlikely to be taken at face value and consumers will be able to critically assess veracity claims), artificially generated associations can have long-term harmful effects (see also (Rini and Cohen, 2022)).
Researchers have begun to explore the different perspectives on deepfakes through attitudinal research. Although this body of work is still somewhat nascent, results suggest significant concern about the discovery and propagation of deepfake videos (Gottfried, 2019; Cochran and Napshin, 2021). Additionally, studies have focused on the harms associated with deepfakes, including a study published in 2021 in which respondents deemed creating and posting NSII, even when labeled as fake, as highly harmful and deserving of punishment (Kugler and Pace, 2021). One study (Fido et al., 2022) found that while female respondents indicated greater levels of victim harm, men were more likely to create NSII. The study also found that harms are greater when content is shared, as opposed to when it is used for personal sexual gratification. To the best of our knowledge, a study with Indian respondents is the sole general population survey outside of the United States. They found that awareness of deepfakes among Indian respondents was low, and even those who knew about deepfakes expected (incorrectly) to be able to easily discern fake videos (Shahid et al., 2022).
A subset of research has also been conducted with respondents who are knowledgeable about deepfakes. These studies have focused primarily on ethical or privacy concerns. One study examined Reddit users actively discussing deepfakes, which found, via content analysis, that users were generally pro-deepfake technology, regardless of the consequences (Gamage et al., 2022). In contrast, qualitative research involving individuals using an open-source deepfake creation tool found that participants had reflexively engaged with the concepts of consent and privacy, and expressed significant concerns about the misuse of deepfakes (Widder et al., 2022). These themes were echoed and expanded upon in a mixed-methods study carried out in China (Li and Wan, 2023), where users familiar with deepfakes highlighted informed consent, privacy protection, and non-deception as relevant factors influencing the social acceptance of deepfakes. That is, if deepfakes are transparent and created or distributed with the knowledge and consent of those featured, they are more acceptable than deepfakes created non-consensually and/or for deceptive purposes. In addition to these factors, attitudes are also informed by the content or aim of the deepfakes (e.g., for creative as opposed to malicious purposes). For example, two separate studies found that entertainment (Li and Wan, 2023) or humorous/amusement value (Cochran and Napshin, 2021) mitigated ethical judgments on deepfakes. This may suggest that content such as NSII or parody videos could be considered more acceptable than videos intended to further political disinformation, for example.
In sum, there is some quantitative, survey-based research suggesting that average respondents have concerns about the deleterious potential of deepfakes. Similarly, existing qualitative research with participants who have expertise on deepfakes shows that they share concerns about risks and harms. More research, however, is needed to investigate geographic diversity in general attitudinal quantitative research, as well as prevalence data, to broadly understand deepfake behavior. More work is also needed specifically on attitudes towards creating, sharing, or explicitly seeking out NSII.
2.2. Deepfake Creation and Detection
While researchers and journalists have been documenting the phenomenon of deepfakes for several years, interest has accelerated in recent years due to technological advancements that facilitate the creation of NSII in affordable and low-friction ways (Bond, 2023). For example, ”deepfake pornography,” as a keyword search within Google Scholar, returns 284 results from 2017-2019 as compared to 1,480 results from 2022-2023. Much of the research in this space (Gamage et al., 2021) focuses on documenting how this technology works or highlighting the performance issues for various detection tools or methods (Pu et al., 2021; Dolhansky et al., 2020; Masood et al., 2023; Stroebel et al., 2023; Rana and Sung, 2023). Like other general artificial intelligence research focusing on the issues associated with biased training data sets, deepfake detection research has identified diversity and bias problems within relevant datasets (Xu et al., 2022; Trinh and Liu, 2021). For example, detection techniques have been shown to be sensitive to gender, performing worse on female-based deepfakes (Nadimpalli and Rattani, 2022) as compared to male counterparts. A recent systematic review of deepfake detection papers succinctly concludes that current models are not robust to either novel or challenging conditions (Stroebel et al., 2023). This is particularly relevant to NSII, as non-digitally altered image-based sexual abuse material may be captured surreptitiously, or made with poor cameras and/or bad lighting conditions. NSII is thus likely to demonstrate the exact risk factors for poor detection rates.
In addition to technical tool failures, studies have documented the fallibility of humans in correctly identifying deepfakes as such (Köbis et al., 2021; Korshunov and Marcel, 2020; Thaw et al., 2020). This is seen across a variety of types of deepfakes, from deepfake audio (Mai et al., 2023) to images (Bray et al., 2023; Nightingale and Farid, 2022) and videos (Groh et al., 2022). It is possible that as deepfakes become more ubiquitous, humans will make improvements in identifying markers, however from a technological detection perspective, it seems likely that an arms race of sorts will develop, such that the increasing sophistication of deepfakes will require detection methods to make significant improvements to catch up and keep pace.
2.3. Deepfakes and the Law
The societal implications of deepfakes technology are far-reaching and, while many have been recognized or anticipated, current legislative remedies for misuse are typically deemed unlikely to be effective by experts (Chesney and Citron, 2019; Gieseke, 2020; Delfino, 2019; Mania, 2020; Solichah et al., 2023). Although many scholars call for legislation, there are others who are skeptical of criminalizing NSII or deepfakes more broadly (Kirchengast, 2020; Feeney, 2021; Widder et al., 2022). Insufficient NSII laws should be considered the norm, rather than the exception, when considering the history of failed legislative proposals for image-based sexual abuse more broadly (Carter, 2022; Walker and Sleath, 2017; Akerley, 2020). In the legal sphere, the first American federal law regarding deepfakes was signed in December, 2019. This law focused on the potential of deepfakes to be weaponized by foreign powers to influence elections, and it also established a competition to encourage research into detection technologies (Ferraro et al., 2020). Four states (California, Virginia, New York, and Georgia) have also passed laws banning NSII specifically (Mania, 2020; Ryan-White et al., 2022; Kugler and Pace, 2021). At the end of 2022, a federal bill was introduced in the U.S. House of Representatives specifically prohibiting NSII (Morelle, 2023). Although legislative remedies have approached political deepfakes and NSII separately, one critique from scholars is that both operate in similar ways to silence critical speech (Maddocks, 2020).
Outside of the United States, several countries have either explicitly included NSII in existing laws or are considering making legislative amendments. For example, in Australia, some states and territories make it clear that the non-consensual distribution of intimate images includes digitally altered material as a criminal offense. The Australian Online Safety Act 2021 (of the eSafety Commissioner Australia, 2023) also provides a civil remedy scheme for a range of online harms, including the nonconsensual sharing of digitally altered images. South Korea specifically prohibits NSII, alongside all other forms of pornography (Lee, 2021) and in the England and Wales, the UK Online Safety Act 2023 criminalizes threatening or sharing intimate images without consent, including images that have been digitally altered (Lawson, 2023; eSafety Commissioner, nd). In most countries, however, and including most of those surveyed in our study (Belgium, Denmark, France, Mexico, Netherlands, Poland, Spain), there are no protections at all.
3. Methods
3.1. Ethical Considerations
In mid-2023, we conducted a cross-country online survey on image-based sexual abuse, defined as the non-consensual taking, creating, or sharing of intimate images (photos or videos), including threats to share intimate images. The study procedure and survey were approved by the Human Research Ethics Committee at RMIT University. Respondents were informed of the topic at the beginning of the survey and were allowed to opt out at any time. All questions involving self-reporting of victimization and/or perpetration were voluntary and respondents could choose an answer of ”prefer not to say” while still completing the survey. Respondents were given support service information tailored to their geographic location at the end of the survey. Respondents were compensated based on pilot tests of time taken in each country and exchange rates.
3.2. Study Design
We conducted the online survey in ten countries, employing a research and survey firm (YouGov) to survey a minimum of 1,600 adults (18+) in each country, for a total respondent count of 16,693. YouGov was selected due to their comprehensive country coverage and large panels (Sandle, 2021), which should reduce the likelihood of needing extensive within-country weighting. YouGov retains demographic information on their survey panelists, and screener questions were then used to achieve representative sampling based on quotas. Upon qualifying for the survey, completion rates ranged from 71.0% (Belgium) to 89% (Australia). The final dataset consisted of respondents who passed the YouGov standard quality checks, including bot catchers, speeder detection, and manual checking of open-text questions. In conducting our survey, we intentionally chose a variety of countries which represent the spectrum of how legislation has dealt with NSII, image-based abuse, and pornography more broadly. The goal was to obtain representative samples by age, gender, and location (e.g., state or territory in Australia, province in Korea, etc) within each country, using the latest official population estimates reflecting each surveyed country. The survey built on an earlier survey conducted in Australia (Powell and Henry, 2019), which focused on non-consensual intimate imagery more broadly (but did not specifically ask about deepfakes). The revised survey (median completion time 18.2 minutes) added some additional scales, teased apart how content was obtained (e.g., filmed/photographed versus stolen from a device/cloud), and included questions on NSII specifically.
In this paper, we focus on the subset of 23 questions in our survey relating to NSII and digitally altered intimate images. To administer the survey in the official language(s) of each country, the survey vendor provided translations into French, Spanish, Danish, Dutch, Polish, and South Korean. These were then double-checked and edited as needed by native speaker colleagues of the research team for quality assurance. We aimed to have as similar language as possible across countries, while also accounting for cultural differences in comfort speaking about sexual topics. The survey was deployed from May to June 2023. In some countries, samples were not fully representative. Accordingly, when presenting dis-aggregated findings, the data are weighted by age, gender, and location. When we aggregate across countries, the data are additionally weighted by population, rounded to the nearest 1,000. The weighted statistics are presented throught the quantitative section, and all quantitative analyses were performed using R Statistical Software (v4.3.2; R Core Team 2023)(R Core Team, 2023). Statistics presented in the qualitative section are unweighted. We primarily used the survey (Lumley, 2023), srvyr (Freedman Ellis and Schneider, 2023), marginaleffects (arel2023marginal), and epitools (Aragon, 2020) packages for statistical analyses. When calculating confidence intervals for proportions, we use the Rao-Scott scaled chi-squared distribution for the loglikelihood from a binomial distribution (e.g., % of respondents who had experienced victimization). When calculating within-subject mean differences, we use two tailed one-sample weighted t-tests with bootstrapped standard errors, using 1,000 replicates (e.g., differences in criminalization attitudes across different types of behaviors). When calculating between-gender mean differences we use two tailed two-sample weighted Welch t-tests using bootstrapped standard errors, using 1,000 replicates (e.g., difference in how much women think sharing deepfake pornography of celebrities should be criminalized as compared to men). We used the svyglm function with a quasibinomial family to calculate risk ratios for binary categorical values (e.g. comparing how much more likely men were to report having watched deepfake pornography as compared to women). When evaluating differences amongst categorical variables in the unweighted qualitative data, we use Pearson’s Chi-squared test with Yates’ continuity correction.
The mean respondent age was 46.0 years ($sd=16.7$). Women accounted for 50.9% of the respondents, men for 47.6%, and together ”other,” ”prefer not to say,” and ”non-binary” accounted for the remaining 1.4%. Detailed demographic breakdowns by country are available in Table A1 in the Appendix. Our gender results and statistical tests focus primarily on the users who identified as women or men, excluding the 1.0% of other users for two reasons: we are using census weights, which are often limited to binary genders, and our total sample (unweighted) of non-binary respondents was 93, which limits our ability to perform statistical tests for significance.
3.3. Measures and Analysis
In relation to NSII, all respondents were asked three closed-ended questions, as well as one optional open-ended question presented towards the end of the survey. They were first asked whether they were familiar with the concept of deepfake pornography using the following wording: ”Realistic-looking fake porn can now be created using artificial intelligence (AI) to swap the faces of pornographic actors with other people’s faces, so that it looks like they’re in a porn video. This is sometimes referred to as ’deepfake pornography’ as it looks very realistic and can be hard to recognize as fake or digitally created. How familiar are you with this concept?”. This question offered four possible categorical response options, as seen in Table 1. Second, respondents were asked how much they thought a series of behaviors were worthy of criminalization (behaviors can be seen in Figure 1). Response options ranged from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with the midpoint of 0 being ”Not sure.” And third, respondents were asked how much they agreed or disagreed with the following statement: ”People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ‘deepfake’) of them without their permission.” Respondents answered this question on a 7-point Likert scale, anchored at 1=Strongly disagree and 7=Strongly agree (we present the mean alongside the median in the results - see below) (Sullivan and Artino Jr, 2013).
The remaining close-ended questions were asked of a subset of respondents, depending on their knowledge and experiences. A subset of respondents who indicated at least some familiarity (i.e., ”I know a little bit about it” or ”I am quite familiar with this”) were asked follow-up questions around NSII behaviors as seen in Figure 2. As part of a broader set of questions regarding the victimization and perpetration of image-based sexual abuse since the age of 18, we also asked the subset of respondents who indicated a previous experience with digitally altered images whether the relevant content was a video, and if so, whether the video was a deepfake (allowing respondents to select yes/no/don’t know, or prefer not to answer). For victimization questions, the wording was intentionally not blame-casting. Similarly, perpetration questions were phrased neutrally to focus on whether the respondent had ever done an action.
All respondents could opt into answering the open-ended question with the following wording: ”Would you like to add any additional thoughts on the topic of deepfakes? Please feel free to skip this question if you don’t have anything to add.” We analyzed the qualitative data for the open-ended response questions using an iterative, inductive approach. Our approach was to translate the responses from origin language to English via the Google Translate API , familiarize ourselves with the data, develop a codebook, code the data, and then review the data to make sense of the themes from the first coding process. To do so, two members of the research team read through 25% of the sample responses across the countries and then co-developed a codebook that detailed descriptions and restrictions on what could be included within a code and provided concrete examples of each code. The same two members of the research team coded the first 150 responses based on four preliminary codes that emerged from the initial review of the data. Coder 1 and coder 2 independently moved through the subset of data labelling the open-ended responses with relevant codes. Following this, coder 1 and 2 compared coding allocations for quality assurance. The two coders then identified and removed responses that were deemed inappropriate or irrelevant to the research (e.g., ”no comment,” or ”no thanks, I do not know this topic”). Finally, the team revised and discussed the themes that emerged from the responses (n = 2,474). The final codebook contained four key themes: ”condemnation,” ”punitive and justice attitudes,” ”a fear of technology,” and ”victim-blaming.” Interrater reliability (IRR) for each code was calculated using Cohen’s Kappa. Across the four codes, Kappa values ranged from .135 to .225, with a median of .173.
4. Results
Results are organized into four sections. The first section addresses the first research question, focusing on awareness and attitude questions that were asked of all respondents. The second section answers the second question by focusing on behaviors involving NSII, including victimization, perpetration, and consumption. The third section investigates the role of gender in influencing answers to the first and second research questions. Finally, the last section is a qualitative analysis of the open-ended opinion question. We present our results primarily descriptively and holistically, particularly at the country level, as we have no reason to a priori assume differences between countries, given the lack of prior data. When discussing gender differences, statistical analyses (accounting for weights) are presented, alongside measures of uncertainty. We conclude with some qualitative trends from the single, open-end question to which respondents could choose to respond.
4.1. Awareness and Attitudes
Table 1. Respondent Familiarity with Deepfakes, by Country (weighted)
Country Response % of Respondents 95% CI
Australia I have never heard of this until now 36.8 35.4-38.2
I have only heard the phrase 30.3 29-31.7
I know a little bit about it 24.7 23.4-26
I am quite familiar with this 8.2 7.3-9.1
Belgium I have never heard of this until now 40.1 39.1-41.1
I have only heard the phrase 28.0 27.1-28.9
I know a little bit about it 25.6 24.7-26.5
I am quite familiar with this 6.3 5.8-6.8
Denmark I have never heard of this until now 34.8 34.1-35.5
I have only heard the phrase 44.1 43.4-44.8
I know a little bit about it 17.0 16.5-17.6
I am quite familiar with this 4.10 3.8-4.4
France I have never heard of this until now 55.2 52.8-57.6
I have only heard the phrase 21.6 19.7-23.6
I know a little bit about it 18.4 16.6-20.3
I am quite familiar with this 4.8 3.8-5.9
Mexico I have never heard of this until now 51.7 48.8-54.7
I have only heard the phrase 22.3 19.9-24.8
I know a little bit about it 21.0 18.7-23.5
I am quite familiar with this 4.9 3.8-6.3
Netherlands I have never heard of this until now 29.0 27.9-30.1
I have only heard the phrase 31.7 30.6-32.8
I know a little bit about it 31.8 30.7-32.9
I am quite familiar with this 7.5 6.9-8.2
Poland I have never heard of this until now 50.9 49.1-52.8
I have only heard the phrase 30.1 28.4-31.8
I know a little bit about it 15.2 13.9-16.5
I am quite familiar with this 3.8 3.2-4.6
South Korea I have never heard of this until now 31.1 29.1-33.2
I have only heard the phrase 39.2 37.1-41.4
I know a little bit about it 26.6 24.7-28.6
I am quite familiar with this 3.0 2.3-3.8
Spain I have never heard of this until now 49.5 47.5-51.5
I have only heard the phrase 27.6 25.8-29.4
I know a little bit about it 19.5 17.9-21.1
I am quite familiar with this 3.5 2.8-4.3
USA I have never heard of this until now 42.5 37.7-47.4
I have only heard the phrase 26.9 22.7-31.4
I know a little bit about it 23.7 19.6-28.1
I am quite familiar with this 6.9 4.5-9.8
On average, 28.1% ($95\%CI=27.0\%-29.3\%$) of respondents reported being either quite familiar with, or knowing at least a little bit about, deepfake pornography. Descriptive statistics including 95% CIs broken out by country are available in Table 1. The percentage of respondents indicating high familiarity ranged from 3.0% in South Korea to 8.2% in Australia.
Figure 1. Heatmap of Mean Criminalization Attitudes by Country. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses
[Heat map with country on the x axis and experience type on the y axis]Heat map with country on the x axis and experience type on the y axis. Mean weighted scores are presented alongside standard errors in parantheses. All scores are positive, indicating that on average, respondents fell on the side of ”should be a crime”.
Figure 1. Heatmap of Mean Criminalization Attitudes by Country. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses
[…]
Finally, respondents were asked about how much they agreed or disagreed with the following statement: ”People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ’deepfake’) of them without their permission,” with anchors of 1=Strongly disagree and 7=Strongly agree. As seen in Table 4.1, across the different countries, the mean response fell between 1 and 2, with a median of 1 in all countries, indicating a general sentiment that people should get upset if they are the target of a deepfake pornography video.
Table 2. Level of agreement with statement: ”People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ‘deepfake’) of them without their permission” (1 = strongly disagree, 4 = neither agree not disagree, 7=strongly agree). Weighted statistics presented.
Country Mean SE Median
Australia 1.8 0.04 1.0
Belgium 1.5 0.03 1.0
Denmark 1.4 0.03 1.0
France 1.7 0.04 1.0
Mexico 1.5 0.03 1.0
Netherlands 1.5 0.03 1.0
Poland 1.6 0.04 1.0
South Korea 1.8 0.04 1.0
Spain 1.4 0.03 1.0
USA 1.8 0.04 1.0
4.2. Behaviors
Behaviors were asked about in three different ways. First, respondents who reported that they were quite familiar with the term ”deepfakes” or knew a little bit about it (28.1%, $95\%CI=27.0\%-29.3\%$) were asked a follow-up question about whether they had engaged in specific deepfake porn behaviors. Second, as part of the larger survey, respondents were asked about general image-based abuse victimization and perpetration. If they reported experiences with digitally altered content, they were asked follow-up questions about whether the digitally altered content were videos, and if so, if the videos were deepfakes. This is distinct from the first question in that 1) it asks for more details of the perpetration/victimization (e.g., whether they had created, threatened to post, or posted such content), and 2) it asks respondents who reported victimization to identify whether they had been the victim of NSII specifically.
4.2.1. Behaviors of ”Familiar” Respondents
The results of the first question (measuring perpetration behaviors) can be seen in Figure 2, where the reported data represents the percentage of all respondents. The respondents who did not see this question due to a lack of familiarity with deepfake pornography are represented in the missing data, which would make each bar sum to 100% if included. For example, 70% of respondents in Spain indicated they had never heard of this until now or they had only heard the phrase, so they were not asked this question. The most common behavior was the viewing of celebrity deepfake pornography, ranging from 2.8% ($95\%CI=2.1\%-3.7\%$) of all respondents in Poland, to 8.0% ($95\%CI=6.6\%-9.5\%$) of all respondents in Australia. Aggregated across the different countries, 6.2% ($95\%CI=5.6\%-6.9\%$) of all respondents indicated that they had viewed celebrity deepfake pornography. The second most common behavior (2.9% of all respondents, $95\%CI=2.5\%-3.3\%$) was the viewing of deepfake pornography involving ”ordinary” people, ranging from 0.7% ($95\%CI=0.4\%-1.3\%$) in Denmark, to 5.1% ($95\%CI=4.1\%-6.2\%$) in Mexico. Behaviors that required more time, skills, or resources were rarer, such as creating or paying someone to create deepfake pornography of ordinary people (0.6%, $95\%CI=0.4\%-0.9\%$, and 0.5%, $95\%CI=0.3\%-0.7\%$), respectively, aggregated across countries)).
Figure 2. Deepfake Porn Behaviors by Country. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey
Faceted bar charts, with % of all respondents on the x axis alongside 95% confidence intervals, and country on the y axis. Rates are quite low for every behavior. The viewing of celebrity and non-celebrity deepfake are the most endorsed behaviors.
Figure 2. Deepfake Porn Behaviors by Country. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey
4.3. Respondents who Reported Personal Victimization or Perpetration
4.3.1. Victimization
Experiences broken out by country are available in Figures 4 and 3. Aggregated across behaviors, respondents from Australia (3.7%), Mexico (2.9%), and South Korea (3.1%) reported the highest rates of victimization by deepfake pornography. Full results with 95%CIs can be seen in Table A2 in the Appendix.
Aggregated across countries, 2.2% ($95\%CI=1.8\%-2.6\%$) of respondents reported some form of victimization. This included: 1.2% ($95\%CI=0.9\%-1.5\%$) of respondents who reported that someone had created NSII content of them; 1.3% ($95\%CI=1.0\%-1.6\%$) who reported that someone had posted or sent NSII content of them; and 1.2% ($95\%CI=0.9\%-1.5\%$) who reported that someone had threatened to post or send NSII content of them.
Figure 3. % of Respondents Reporting Different Types of Deepfake Video Victimization, by Country. 95% CIs in grey.
Faceted bar chart with country on the x axis and % of respondents indicating victimization experiences on the y axis. Victimization experiences include images created, posted, sent, or shown without consent, or threats made to post, send, or show images. Rates are lowest in Belgium across behaviors, and highest in Australia, and fall between 0.3% and 2.6%
Figure 3. % of Respondents Reporting Different Types of Deepfake Video Victimization, by Country. 95% CIs in grey.
4.3.2. Perpetration
Aggregating across behavior types (creating, posting, threatening to post), self-reported NSII perpetration was rare (1.8%, $95\%CI=1.4\%-2.2\%$), with the highest rates reported in Australia (2.4%, $95\%CI=1.7\%-3.4\%$) and the United States (2.6%, $95\%CI=1.9\%-3.5\%$), and lowest in Belgium (0.3%, $95\%CI=-0.1\%-0.7\%$). Full results can be seen in Table A2 in the Appendix. Aggregating across countries, 1.0% ($95\%CI=0.8\%-1.4\%$) of respondents indicated creating deepfake pornography, 1.0% ($95\%CI=0.8\%-1.3\%$) reported threatening to post, send, or share deepfake pornography, and 0.7% ($95\%CI=0.5\%-0.9\%$) reported actually posting, sending, or sharing deepfake pornography content. We included the creation of deepfake pornography as a perpetration behavior to mirror the fact that ”creation” behaviors, such as filming or photographing real content without consent, fall under the umbrella of image-based abuse perpetration. Nevertheless, we acknowledge that such behaviors might not be considered as ”perpetration” by others.
Figure 4. % of Respondents Reporting Different Types of Deepfake Video Perpetration, by Country. 95% CIs in grey.
Faceted bar chart with country on the x axis and % of respondents indicating perpetration behaviors on the y axis. Perpetration behaviors include images created, posted, sent, or shown without consent, or threats made to post, send, or show images. Rates are low in Belgium, Poland and the Netherlands, and highest in Australia and the United States and fall between 0.1% and 1.6%
Figure 4. % of Respondents Reporting Different Types of Deepfake Video Perpetration, by Country. 95% CIs in grey.
4.4. Gender
4.4.1. Awareness and Attitudes
[…]
Figure 5. Heatmap of Mean Criminalization Attitudes by Respondent Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parantheses.
Heatmap showing attitudes towards various image-based sexual abuse behaviors. The x axis is gender (women and men), and the y axis is each type of behavior (e.g., posting, sending, or showing real intimate content without consent). The behaviors involving real content all fall between ”should be a crime” and ”definitely should be a crime,” and all weighted means are ¿0, indicating that all of the behaviors fell between ”not sure” and ”definitely should be a crime.”
Figure 5. Heatmap of Mean Criminalization Attitudes by Respondent Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parantheses.
[…]
4.4.2. Behaviors
Of the two most commonly reported behaviors, men respondents were more likely to report viewing deepfake pornography, both that of celebrities ($F=71.1,ndf=1,ddf=16510,p<0.001,RR=3.5,95\%CI=2.77-4.42$) and non-celebrities ($F=14.3,ndf=1,ddf=16510,p<0.001,RR=2.66,95\%CI=1.91-3.70$).
Figure 6. Deepfake Porn Behaviors by Gender. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
Faceted bar charts, with % of all respondents reporting they have engaged in the behavior on the x axis, and gender on the y axis. Rates were at 1% or less for every behavior, regardless of gender, with the exception of: the viewing of celebrity deepfake pornography (9.9% of men, 2.8% of women), and the viewing of non-celebrity deepfake pornography (4.2% men, 1.6% of women).
Figure 6. Deepfake Porn Behaviors by Gender. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
4.4.3. Respondents who Reported Personal Victimization or Perpetration
Looking at the effects of gender on victimization and perpetration, men reported significantly higher rates across all victimization and perpetration experiences related to creation and threats. For example, the risk of men reporting having been threatened with deepfake dissemination was 2.6 times that of women, and the risk of them threatening someone else was 1.9 times that of women. Full results are presented below in Table 4.4.3. Notably, there were no significant differences by gender in the risk of disseminating content or having their content disseminated.
Table 3. Experiences of NSII by Gender, with Risk Ratios.
Type Experience Proportion of Proportion of p-value RR 95%CI
”yes” men ”yes” women
Perpetration
Creating 0.015 0.006 ¡.005 2.31 1.29-4.15
Threatening 0.015 0.006 ¡.01 2.64 1.40-4.98
Posting/Sharing 0.009 0.005 0.16 1.63 0.82-3.25
Victimization
Created 0.016 0.007 ¡.001 2.230 1.39-3.58
Threatened 0.016 0.008 ¡.001 1.91 1.18-3.10
Shared 0.015 0.010 0.093 1.52 0.93-2.46
4.5. Qualitative Responses
In the survey, respondents were asked an optional, open-ended question: ”Would you like to add any additional thoughts on the topic of ‘deepfakes’? Please feel free to skip this question if you don’t have anything to add.” A total of 2,477 open-ended responses were collected. Of those responses, 1,180 were removed due to inappropriate or irrelevant responses. This left us with 1,296 open-ended responses (7.8% of respondents). No respondents discussed their own experiences of either victimization or perpetration.
[…]
The second most prevalent theme, found in just over 30% of all responses to the open-ended question, referred to a fear of technology; specifically deepfakes and artificial intelligence (AI). There were significant differences by country ($x^{2}=84.89,df=9,p<0.001$). These responses were the highest across Australia (22.9%), the United States (18.3%), and Mexico (13.4%), and the lowest in France (2.9%) and Belgium (4.4%). References to the futuristic nature of deepfake technology, fear of its capabilities and accessibility, as well as the realness of deepfaked images, showed up across these responses. For example, one man from Mexico (aged 25-34) commented on how “it is difficult to assimilate [to] these types of tools, before they seemed futuristic but they are already very normal.” Many respondents described deepfakes as a troubling and terrifying advancement of AI technology. Some, particularly women, indicated their concern about being victimized by deepfake pornography. For example, a woman from Denmark (aged 25-34) referred to deepfakes as “widely scary” because “it’s hard not to think if you have become a victim yourself without knowing it.” Other respondents focused on the continued growth of deepfakes and the implications they might have for future generations. For example, an Australian woman (aged 45-54) described how she “fear[ed] for [her] daughter’s future.” Comments on the realness of deepfakes was a common thread throughout responses, with a Polish man (aged 35-44) saying “soon it will not be known if what we [are] seeing is reality or not.” Similarly, a Belgian woman (aged 45-54) questioned “can we still know if the photo or video seen or received is real or not?! It becomes difficult to make your own judgement.” Respondents also spoke to the difficulties of regulating deepfakes online. A man from Mexico (aged 65+) emphasized how it is “almost impossible to control,” while a woman from the United States (aged 65+) asked “how can one protect oneself from this? Or prove it’s not them?”
The third most common theme was punitive and justice attitudes. Across responses, men were slightly more likely to express punitive and justice attitudes ($X^{2}=4.85,df=1,p<0.05$). References to punishment, injustice, criminalization, and governmental responsibility were common in South Korea. Just over half (55%) of South Korean respondents demanded strengthened laws relating to deepfakes. Responses often referred to deepfake pornography as a “sex crime” and called for the implementation of strong criminal justice regulations and sanctions. Accordingly, many South Korean respondents argued for the development of criminal sanctions that imposed “strong punishment” for perpetrators of deepfake pornography. As one South Korean respondent (male, aged 45-54) stated, “deepfakes [are] not yet controversial, but it can cause many social problems. In reality, legal regulations are needed to strongly punish [these] sex crimes.”
Many respondents disclosed that they were unaware if penalties and sanctions were available in their country for deepfakes, while others were concerned with the slow pace of governmental responses. One woman from the United States (aged 55-64) stated that she “feared that politics doesn’t attract enough experts in tech who could help shape policy and procedure on deepfakes … and it needs to be something for which governments set as a goal to get ahead of and not ignore [or] play catch up.” Similarly, a woman from Denmark (aged 25-34) emphasized the need to “hurry to regulate how people are taking advantage of AI.” References to internet regulation and platform responsibility were found across some responses. For instance, one Australian woman (aged 65+) commented that “platforms need to take more responsibility for content,” and a man from the Netherlands (aged 65+) recommended that this type of content should be “immediately removed by relevant platforms and should be punishable for both maker and distributor.”
Finally, only about 4% of responses indicated victim-blaming attitudes. These respondents noted that people should be more careful with what they post on the internet, and that they should refrain from taking or sharing nude photos or videos, and stop complaining about the issue and take responsibility. These responses were highest in Australia (32.0%) and lowest in South Korea (0.0%). For example, a non-binary respondent from Mexico (aged 35-44) and a woman from Poland (aged 55-64) noted that people have to be careful with what they share on the internet. Two Australian respondents encouraged people to stop taking nude images altogether; for example, one Australian man (aged 55-64) commented “if you don’t want something shared on the internet don’t take nude photos” while an Australian woman (aged 65+) said, “that’s why I do NOT send photos over the internet at all. I can’t understand why people send photos to anyone — keep your [images] private.” One man from Poland (aged 65+) was more explicit and indicated that people should not complain about being victimized: “you just should not share any photos on the web at all, and if you make it available, you should take into account the consequences and then regret pretending to be a fool and even worse to blame anyone.”
In summary, only 7.8% of all respondents provided relevant answers to the optional, open-ended question. This means that the majority of survey respondents did not provide additional thoughts on the topic - either because they did not know enough about deepfake pornography or for other reasons.
5. Discussion
Our study generated three main findings about NSII awareness, behaviors, and gender dynamics. First, despite widespread press coverage, particularly in Western countries, the concept of deepfake pornography is not universally known, yet on balance when informed about the concept, respondents thought behaviors associated with non-consensual fake intimate imagery should be criminalized. Second, the most common behaviors self-reported by respondents were passive and readily available (e.g., the consumption of celebrity deepfake pornography), with behaviors requiring effort and/or money being very rare (e.g., creating deepfake pornography). Third, compared to women, men think of these behaviors as less bad, and are more likely to also report being both victimized by and perpetrating NSII.
5.1. Attitudes and Awareness
Less than 30% of all respondents indicated some level of familiarity with the concept of deepfake pornography. This was also arguably reflected in the qualitative responses, with only 7.8% of all respondents answering the open-ended question. Presumably, as the technology becomes more widespread and ever more available, familiarity will increase. This may be a double-edged sword; it is likely to result in increased use of NSII as an abuse vector (both for purposes of humiliation, and as a tool for sexual extortion), but may also result in increased suspicion of the veracity of explicit content. One positive finding was general agreement across countries that perpetration behaviors associated with ”traditional” image-based sexual abuse are deserving of criminalization. Research has widely found some degree of victim-blaming reflected in attitudes towards traditional non-consensual intimate imagery (Flynn et al., 2023; Mortreux et al., 2019; Sugiura and Smith, 2020). Attitudes were slightly more lenient with regard to deepfake behaviors, although over time attitudes may evolve with increased familiarity. For instance, the realization that any private individual can be deepfaked, requiring no actions on the part of the target, may engender more empathy and less victim-blaming attitudes. It should be noted that few respondents in the qualitative part of the survey expressed victim-blaming attitudes. This can be compared to other more commonly expressed pro-social attitudes in the qualitative responses, including those expressing condemnation or support for punitive action.
5.2. Behaviors
The vast majority of respondents indicated no engagement in any type of NSII behaviors, with the most popular behavior being the consumption of celebrity deepfake pornography. It was even rarer for respondents to indicate perpetration themselves. This was similar for victimization, with few respondents indicating that they had experienced deepfake pornography. However, the victimization statistics in particular should be interpreted with caution. Even with traditional forms of image-based sexual abuse, many victims may be unaware of their victimhood. NSII may exacerbate this dynamic given that deepfake technology ”democratizes” victimization and image-based sexual abuse by enabling anyone with a computer to be a perpetrator, and anyone with imagery online to be targeted. Also, there are additional unique motivations for NSII that may result in less likelihood of victim notification; for instance, if the motivations for the creation or distribution of NSII are for sexual gratification purposes and/or for demonstrating or practicing digital skills in AI. Finally, as content may be distributed on dedicated sites or within closed communities on sites such as Discord or Reddit, victims of NSII may not know about or find out about their images.
As a subset of image-based sexual abuse, it is useful to compare rates of NSII with previously reported statistics on image-based sexual abuse more broadly. There are several associated challenges, including sampling methodologies and definitions. For example, studies may define image-based sexual abuse behaviors differently, failing to include the non-consensual creation or filming of content, or threatening to share content, alongside the actual non-consensual sharing of content. Studies may also intentionally survey younger populations who may be more likely to have relevant experiences, as opposed to collecting population-level prevalence data. Notwithstanding those limitations, previous work has shown that image-based abuse prevalence is increasing over time, driven by changes in technology use, dating, sex, and privacy (Powell et al., 2020). Powell et al. (Powell et al., 2020) found on average, across Australia, New Zealand, and the UK, that  1/3 of survey respondents indicated at least one form of image-based sexual abuse, defined similarly to our study. An American-based study (Ruvalcaba and Eaton, 2020) focused exclusively on the non-consensual sharing of content, and found that 1/12 respondents indicated victimization, while 1/20 reported perpetration. Even taking the more expansive definition of image-based sexual abuse, our rates of perpetration and victimization using deepfake technology were quite low. Given the trends in traditional image-based sexual abuse (Powell et al., 2020), and the increasing ease of NSII creation, we expect these rates to increase over time.
5.3. Gender and NSII
Awareness of NSII was higher among respondents who identified as men as compared to women. This is consistent with the literature on the ”digital divide” which suggests that women face disproportionate barriers to access and use to the internet compared to men. Even when women have access to the internet and technology, they may not have the time, resources, skills, or knowledge to leverage it (Acilar and Sæbø, 2023). Additionally, the attitudinal questions indicate a trend towards men diminishing the harms or ”badness” of NSII behaviors. This is consistent with previous findings on NSII (Fido et al., 2022), and with other findings around image-based sexual abuse more broadly (Attrill-Smith et al., 2021; Bothamley and Tully, 2017).
Men were more likely to report both perpetration and victimization by NSII, across subtypes. Studies continue to report mixed findings regarding the role of gender in image-based sexual abuse. Powell et al. (Powell et al., 2020) reported higher rates of perpetration in men, and similar rates of victimization for men and women, but differential impacts. Our findings are somewhat inconsistent with a report which found the majority of NSII available online targeted women (Adjer et al., 2019). We argue that the differences in sampling methodology may help explain these inconsistencies. For instance, in the event of sexual extortion, the content may not be made available online. It is possible that the trend seen in gender may be somewhat explained by the motivation of the perpetrator, given previous research to support higher rates of ”sextortion” in men as compared to heterosexual women (Eaton et al., 2023), including for adolescent populations (Patchin and Hinduja, 2020). Another explanation for this discrepancy is that research suggests most deepfake videos online (an estimated 96%) are of women actors and musicians, who may be less likely to participate in survey research than non-public figures like our survey respondents (Adjer et al., 2019).
While we asked broadly about motivations in the larger survey, we did not ask specifically ask about motivations for NSII perpetration or victimization. Motivations for NSII behaviors may range from personal sexual gratification, to use for extortion purposes, to demonstrating technical skill in the creation of content. A consideration for future research should be the use of NSII for extortion, particularly given anecdotal evidence of significant associated harms (e.g., suicide, suicidal ideation, or suicidal attempts) (Pu et al., 2021).
5.4. Limitations and Mitigation
Cross-country surveys are logistically difficult and require significant consideration in interpreting findings. Our study had to address language differences and non-response bias. The former were addressed, as detailed above, through cross-validation. The latter was a more significant concern, as others have noted that language used to discuss a topic such as image-based sexual abuse and pornography more broadly will be culturally dependent (Schoenebeck et al., 2023). We attempted to mitigate this by allowing respondents to select ”Prefer not to say” for any question that involved self-reporting on behavior, as well as periodically reminding respondents of that option throughout the survey. Nevertheless, it is possible that selection bias may have resulted in an underestimation of prevalence, with respondents opting out of the survey due to personal experiences or demographic characteristics. South Korea, for example, had a high percentage of bisexual respondents. This may indicate that our attempts at mitigating non-response bias were less successful in South Korea. We note that, despite explicit goals for representative sampling, the participant pool fell short, necessitating the need to weight results.
5.5. Design Considerations
In this study, we found that the majority of respondents across all countries were relatively unfamiliar with the concept of deepfake pornography. We also found that respondents, on average, found deepfake pornography behaviors to be deserving of punishment. Reported perpetration and victimization rates were low, with the most common behavior reported by respondents being the consumption of deepfake pornography. Finally, gender did play a part in both experiences and attitudes, with men deeming deepfake behaviors as less harmful or deserving of punishment as compared to women. This has implications for both policy and practice.
As relatively few respondents were aware of deepfake pornography, this signals first and foremost that more education is needed to both deter potential perpetration and/or the consumption of NSII. Education efforts should focus on the harms suffered by victim-survivors, as well as the potential consequences of perpetration. Given the gender dynamics in terms of perceived harmfulness, prioritization around education should focus on reaching boys and men.
Second, many scholars have focused on the need for legislation to criminalize these image-based abuse behaviors and provide clear redress for victims (Chesney and Citron, 2019; Gieseke, 2020; Delfino, 2019; Mania, 2020; Solichah et al., 2023). In order for a law to have both general and specific deterrent effects, the consequences of breaking that law has to be known by the general public, and someone who does break the law has to be prosecuted and appropriately punished. Some of the trends, particularly the high rates of prevalence in both South Korea and Australia, indicate that laws are not a panacea to eliminate or even reduce this behavior. For instance, Australia criminalizes the distribution of non-consensual intimate imagery, including NSII (Australia, 2021), while South Korea has laws allowing for up to 3 years imprisonment for merely viewing non-consensual intimate imagery (Center, 2019; Maeng and Lee, 2022). Given that this behavior primarily happens online, that victim-survivors may not be aware of the material, or if they become aware, may be reluctant to report it, it can be argued that existing legislation provides weak deterrence since very few people have been prosecuted for NSII in jurisdictions that do have such laws in place. As we found in this study, there was also low awareness of the concept of deepfake pornography, and one may reasonably infer low awareness of related laws. While developing legislation around an issue helps create societal norms and signals the unacceptability of a behavior, it is crucial to concede the limitations of passing regulations in enacting true societal change. This also allows us to develop parallel efforts, such as technical solutions, that may mitigate these harms by reducing the creation and dissemination of NSII.
Third, there are a number of points of potential intervention for technical solutions. First, discovery tools that allow individuals to search the web for sexual material via facial recognition technology may be empowering, although there are safety, privacy, and security concerns that need to be taken into account (e.g., building in protections such that individuals can only search for content of their own faces). Second, while many videos are posted on dedicated deepfake websites, or make no claims as to veracity, the use of deepfakes for sexual extortion indicate that technical methods of annotating content, such as watermarks, may help mitigate NSII harms, although not fully (Harris, 2021). Third, platforms can make choices around what they allow and disallow in their policies and community standards (legislation can assist in justifying the choices made by platforms). Google Search (goo, 2023), Reddit (Hollister, 2023), and PornHub(Cole, 2018) have already published dedicated NSII policies. Fourth, various stakeholders have worked to help create solutions for victims, primarily focusing on triaging after victimization has occurred. It will be necessary for these help-seeking resources (whether helplines, online forms, or chatbots(Falduti and Tessaris, 2022; Maeng and Lee, 2022)) to be aware of this new type of image-based sexual abuse and up-to-date with the types of redress that victims may seek (Aliman et al., 2021). Fifth, where tools do have a payment or sign up flow, users could be required as part of the flow to ingest and acknowledge information about the harms and potential consequences of NSII creation and dissemination. Finally, platforms have demonstrated an ability to detect nude content or, in some cases, non-consensual nude or sexual content. If creator tools could use those tools to block creation, or at the very least serve a warning interstitial about the harms associated with NSII, it may act as a further deterrent for uploading this content.
Figure 7. Potential Intervention Points. The top half of this figure represents societal-level interventions, while the bottom half is more directed at individuals or companies in the deepfake econsystem.
Temporal ordering of potential intervention points, starting at the pre-creation phase (highlighting UX interventions to deter NSII content), moving into the creation phase (highlighting nudity classifiers and watermarking tools), then discovery (potential tooling around ways to discover victimization), and finally recourse (ensuring resources are up to date and useful for NSII experiences). The importance of relevant legislation and enforcement overlays all phases.
Figure 7. Potential Intervention Points. The top half of this figure represents societal-level interventions, while the bottom half is more directed at individuals or companies in the deepfake econsystem.
As practitioners and designers think about future considerations for intervention, it may be helpful to break this issue down into relevant behaviors as seen in Figure 7 and consider intervention options (which should be considered complementary, as opposed to solely sufficient) at each point in the process. The increased ease of access to deepfake technologies was the catalyst for this project, and represents the first place at which interventions could be made. Are there ways, for example, to limit these tools’ abilities to create nude or intimate imagery? If creators get to the point where barriers have been hacked, are there warning messages that can either highlight associated harms to targets, or potential punishments associated with creation/distribution of such content? After creation, are there ways to build in indicators of provenance (e.g., watermarking) that signal the synthetic nature of the content? Finally, assuming most impediments can be circumvented by a motivated creator, empathetic and multi-modal resources for victims could include both tooling to discover where relevant content is published, and help for seeking take downs and recourse more broadly. These are primarily design interventions, but overarching all of these would be the societal interventions of educational programming, and relevant legislation making clear potential punishments associated with each type of behavior (extortion, posting/distribution, etc); paired with prosecutorial actions that indicate a willingness to hold perpetrators to account.
5.6. Conclusion
In this study, we used a survey to explore general attitudes toward NSII, as well as NSII behaviors and experiences in ten different countries around the world. We found relatively low rates of awareness across countries, but respondent attitudes (as seen in the questions about whether victims should get upset and criminalization, and in the qualitative responses) were concern and condemnation of image-based sexual abuse behaviors, including NSII. The most popular reported behavior was the consumption of celebrity NSII, while personal victimization and perpetration rates were quite low. Men tended to find consuming NSII less objectionable than women, which was also seen in the higher rates of self-reported consumption behavior. Finally, men were more likely to report both more perpetration and more victimization by NSII, a trend that future research should explore as we may expect prevalence to increase over time. Laws, policies, and support services can help to mitigate the harms of NSII; however, more resources need to be directed toward detecting and blocking the creation of content prior to dissemination, as well as preventing the abuse from happening in the first place through educational programs and initiatives.
Appendix
Figure A3. Heatmap of Median Criminalization Attitudes by Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Numbers shown are medians.
[Heat map with gender on the x axis and experience type on the y axis]Heat map with country on the x axis and experience type on the y axis. Weighted medians are presented. Women and men tend to agree with regard to real content, while women consider the creating, sharing, and viewing of deepfake pornography to be more deserving of criminalization as compared to men
Figure A3. Heatmap of Median Criminalization Attitudes by Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Numbers shown are medians.